# Fast Robust Kernel Regression through Sign Gradient Descent with Early Stopping

### v6

@@ Abstract
Kernel ridge regression, KRR, is a generalization of linear ridge regression that is non-linear in the data, but linear in the model parameters. Here, we introduce an equivalent formulation of the objective function of KRR, which opens up both for replacing the ridge penalty with the $\ell_{\infty}$ and $\ell_{1}$ penalties and for studying kernel ridge regression from the perspective of gradient descent.
Using the $\ell_{\infty}$ and $\ell_{1}$ penalties, we obtain robust and sparse kernel regression, respectively. We further study the similarities between explicitly regularized kernel regression and the solutions obtained by early stopping of iterative gradient-based methods, where we connect $\ell_{\infty}$ regularization to sign gradient descent, $\ell_{1}$ regularization to forward stagewise regression (also known as coordinate descent), and $\ell_{2}$ regularization to gradient descent, and, in the last case, theoretically bound for the differences. We exploit the close relations between $\ell_{\infty}$ regularization and sign gradient descent, and between $\ell_{1}$ regularization and coordinate descent to propose computationally efficient methods for robust and sparse kernel regression.
@@
Keywords: Kernel Ridge Regression, Robust Regression, Sign Gradient Descent, Gradient Descent, Gradient Flow
@@ 1 Introduction
In the linear case, many alternatives to the ridge penalty have been proposed, including the lasso penalty (Tibshirani,, 1996), which is known for creating sparse models. By replacing the ridge penalty of KRR with the lasso penalty Roth, (2004), Guigue et al., (2005) and Feng et al., (2016) have obtained non-linear regression that is sparse in the observations. Rather than using the lasso, or $\ell_{1}$, penalty, Russu et al., (2016) and Demontis et al., (2017) trained kernelized support vector machines with max (or infinity, $\ell_{\infty}$) norm regularization to obtain models that are robust against adversarial data, which in the context of regression would translate to outliers.
@@
In this paper, we present an equivalent formulation of KRR, which we use to obtain kernel regression with the $\ell_{\infty}$ and $\ell_{1}$ norms, in addition to the default, $\ell_{2}$, norm. We also use the equivalent formulation to solve kernel regression using the three different gradient-based optimization algorithms gradient descent, sign gradient descent, and coordinate descent (forward stagewise regression). We relate each iterative algorithm to one explicitly norm-based regularized solution and use gradient-based optimization with early stopping to obtain computationally efficient robust, and sparse, kernel regression. Even though robust kernel regression has more evident applications than sparse kernel regression does, since the calculations are analogous for $\ell_{\infty}$ and $\ell_{1}$ regularization we include both cases. However, in the results section, we focus more on the former than the latter.
The rest of the paper is structured as follows. In Section 2, we review kernel ridge regression, before introducing our equivalent formulation in Section 3. In Section 4, we review kernel gradient descent, KGD, and theoretically compare KGD with infinitesimal step size to KRR. In Section 5, we generalize KRR into robust and sparse kernel regression by replacing the $\ell_{2}$ penalty with the $\ell_{\infty}$ and $\ell_{1}$ penalties, respectively. Similarly to in Section 4, we also introduce gradient-based optimization algorithms that we relate to the explicitly regularized methods; kernel sign gradient descent for robust kernel regression, and kernel coordinate descent for sparse kernel regression. Finally, in Section 6, we demonstrate our findings with experiments on real and synthetic data.
@@
All proofs are deferred to Appendix C.
@@ 2 Review of Kernel Ridge Regression
For a positive semi-definite kernel function, $k(\bm{x},\bm{x^{\prime}})\in\mathbb{R}$, and $n$ paired observations, $(\bm{x_{i}},y_{i})_{i=1}^{n}\in\mathbb{R}^{p}\times\mathbb{R}$, presented in a design matrix, $\bm{X}=[\bm{x_{1}}^{\top},\bm{x_{2}}^{\top},{\dots}\bm{x_{n}}^{\top}]^{\top}\in\mathbb{R}^{n\times p}$, and a response vector, $\bm{y}\in\mathbb{R}^{n}$, and for a given regularization strength, $\lambda>0$, the objective function of kernel ridge regression, KRR, is given by
@@
Here, $\bm{\hat{f}}=\bm{\hat{f}}(\bm{X})\in\mathbb{R}^{n}$ and $\bm{\hat{f}^{*}}=\bm{\hat{f}^{*}}(\bm{X^{*}},\bm{X})\in\mathbb{R}^{n^{*}}$ denote model predictions for the training data, $\bm{X}$, and new data, $\bm{X^{*}}=[\bm{x^{*}_{1}}{{}^{\top}},\bm{x^{*}_{2}}{{}^{\top}},{\dots}\bm{x^{*}_{n^{*}}}{{}^{\top}}]^{\top}\in\mathbb{R}^{n^{*}\times p}$, and $\bm{K}=\bm{K}(\bm{X})\in\mathbb{R}^{n\times n}$ and $\bm{K^{*}}=\bm{K^{*}}(\bm{X^{*}},\bm{X})\in\mathbb{R}^{n^{*}\times n}$ denote two kernel matrices defined according to $\bm{K}_{ij}=k(\bm{x_{i}},\bm{x_{j}})$ and $\bm{K^{*}}_{ij}=k(\bm{x^{*}_{i}},\bm{x_{j}})$. The weighted norm, $\|\bm{v}\|_{\bm{A}}$, is defined according to $\|\bm{v}\|_{\bm{A}}^{2}=\bm{v}^{\top}\bm{A}\bm{v}$ for any symmetric positive definite matrix $\bm{A}$.
@@
An alternative interpretation of KRR is as linear regression for a non-linear feature expansion of $\bm{X}$. According to Mercer’s Theorem (Mercer,, 1909), every kernel can be written as the inner product of feature expansions of its two arguments: $k(\bm{x},\bm{x^{\prime}})=\bm{\varphi}(\bm{x})^{\top}\bm{\varphi}(\bm{x^{\prime}})$ for $\bm{\varphi}(\bm{x})\in\mathbb{R}^{q}$. Thus, denoting the feature expansions of the design matrix and the new data with $\bm{\Phi}=\bm{\Phi}(\bm{X})\in\mathbb{R}^{n\times q}$ and $\bm{\Phi^{*}}=\bm{\Phi^{*}}(\bm{X^{*}})\in\mathbb{R}^{n^{*}\times q}$, the two kernel matrices can be expressed as $\bm{K}=\bm{\Phi}\bm{\Phi}^{\top}$ and $\bm{K^{*}}=\bm{\Phi^{*}}\bm{\Phi}^{\top}$. Thus, for $\bm{\beta}=\bm{\Phi}^{\top}\bm{\alpha}$, Equations 1 and 3 become
@@ 3 Equivalent Formulations of Kernel Ridge Regression
In this section, we present an equivalent formulation of the objective function in Equation 1, which has the same solution for $\bm{\hat{\alpha}}$. This formulation opens up for generalizing KRR by using penalties other than the ridge penalty and also provides an interesting connection to functional gradient descent (Mason et al.,, 1999).
In Appendix A, we take this one step further by reformulating Equation 1 directly in the model predictions, $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$, by presenting two equivalent objectives, that, when minimized with respect to $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$, generate the solution in Equation 3. In this and the following sections, all calculations are done with respect to $\bm{\alpha}$, and the corresponding expressions for $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ are obtained through multiplication with $[\bm{K}^{\top},\bm{K^{*}}{{}^{\top}}]^{\top}$. However, in Appendix A, we show how the expressions for $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ can alternatively be obtained directly without taking the detour over $\bm{\alpha}$.
In Lemma 1, we show how we can move the weighted norm in Equation 1 from the penalty term to the reconstruction term.
Lemma 1.
@@
Remark 1: Despite the appearance of $\bm{K}^{-1}$ in Equation 5b, this inverse never needs to be calculated, since in all calculations, $\bm{K}^{-1}$ cancels through multiplication by $\bm{K}$.
The alternative formulation in Equation 5b, where the reconstruction term, rather than the regularization term, is weighted by the kernel matrix, has two interesting implications. First, a standard ridge penalty on the regularization term opens up for using other penalties than the $\ell_{2}$ norm, such as the $\ell_{1}$ and $\ell_{\infty}$ norms. Although these penalties have previously been used in the kernel setting, it has been done by replacing $\|\bm{\alpha}\|_{\bm{K}}^{2}$ in Equation 1 by $\|\bm{\alpha}\|_{1}$ or $\|\bm{\alpha}\|_{\infty}$, thus ignoring the impact of $\bm{K}$, by acting as if the objective of KRR were to minimize $\|\bm{y}-\bm{K}\bm{\alpha}\|_{2}^{2}+\lambda\|\bm{\alpha}\|_{2}^{2}$, and thus losing the connection to linear regression in feature space.
@@
$\bm{K}\bm{\alpha}-\bm{y}+\lambda\bm{\alpha}$ (6)
@@
which is the gradient used in functional gradient descent, where it is obtained from differentiating functionals. Here, however, it is a simple consequence of the equivalent objective function of KRR.
@@ 4 Kernel Gradient Descent and Kernel Gradient Flow
In this section, we investigate solving kernel regression iteratively with gradient descent. We also use gradient descent with infinitesimal step size, known as gradient flow, to obtain a closed-form solution which we use for direct comparisons to kernel ridge regression.
The similarities between ridge regression and gradient descent with early stopping are well studied for linear regression (Friedman and Popescu,, 2004; Ali et al.,, 2019; Allerbo et al.,, 2023). When starting at zero, optimization time can be thought of as an inverse penalty, where longer optimization time corresponds to weaker regularization. When applying gradient descent to kernel regression, something we refer to as kernel gradient descent, KGD, we will replace explicit regularization with implicit regularization through early stopping. That is, we will use $\lambda=0$ and consider training time, $t$, as a regularizer.
@@
$\frac{d\bm{\hat{\alpha}}(t)}{dt}=\bm{y}-\bm{K}\bm{\hat{\alpha}}(t),\ \bm{\hat{\alpha}}(0)=\bm{0}$ (9)
whose solution is stated in Lemma 2.
Lemma 2.
The solution to the differential equation in Equation 9 is
$\bm{\hat{\alpha}}(t)=(\bm{I}-\exp(-t\bm{K}))\bm{K}^{-1}\bm{y}=:\bm{\hat{\alpha}}_{\text{KGF}}(t),$ (10)
@@
Remark 1: Note that $(\bm{I}-\exp(-t\bm{K}))\bm{K}^{-1}=\bm{K}^{-1}(\bm{I}-\exp(-t\bm{K}))$ is well-defined even if $\bm{K}$ is singular. The matrix exponential is defined through its Taylor approximation and from $\bm{I}-\exp(-t\bm{K})=t\bm{K}-\frac{1}{2!}t^{2}\bm{K}^{2}+\dots$, a matrix $\bm{K}$ factors out, that cancels $\bm{K}^{-1}$.
Remark 2: It is possible to generalize Lemma 2 to Nesterov accelerated gradient descent with momentum (Nesterov,, 1983; Polyak,, 1964). In this case $\exp(-t\bm{K})$ in Equation 10 generalizes to $\exp\left(-\frac{t}{1-\gamma}\bm{K}\right)$, where $\gamma\in[0,1)$ is the strength of the momentum. See the proof in Appendix C for details.
@@
where we have used Lemma 3.
Lemma 3.
$\left(\bm{K}+\lambda\bm{I}\right)^{-1}=\left(\bm{I}-\left(\bm{I}+1/\lambda\cdot\bm{K}\right)^{-1}\right)\bm{K}^{-1}.$
@@ 4.1 Comparisons between Kernel Ridge Regression and Kernel Gradient Flow
Even though not explicitly stated, all estimates depend on the training data $(\bm{X},\bm{y})$. We further use the notation $\bm{\xi}\sim(\bm{\mu},\bm{\Sigma})$ do denote that the random variable $\bm{\xi}$ follows a distribution such that $\mathbb{E}(\bm{\xi})=\bm{\mu}$ and Cov$(\bm{\xi})=\bm{\Sigma}$.
In Proposition 1, we bound the differences between the KGF and KRR solutions in terms of the non-regularized solutions, $\bm{\alpha_{0}}$, $\bm{\hat{f}}^{0}(\bm{X})$ and $\hat{f}^{0}(\bm{x^{*}})$. In parts (a) and (b), where the difference between the parameter and in-sample prediction vectors are bounded, no further assumptions are needed. In parts (c) and (d), with bounds on individual parameters and predictions, including out-of-sample predictions, some very reasonable assumptions have to be made on the data. For all four bounds, the larger the non-regularized value is, the larger the difference between the KGF and KRR estimates is allowed to be.
Proposition 1.
@@
In Proposition 2, we compare the distances to the observation vector, $\bm{y}$, of the KGF and KRR solutions. According to parts (a) and (c), for a given regularization, the furthest possible (among all values of $\bm{y}$, or expected $y_{i}^{2}$) normalized KGF solution lies closer to the observations than the furthest possible normalized KRR solution does. Analogously, according to parts (b) and (d), for a given regularization, the furthest possible KRR solution lies closer to zero than the furthest possible KGF solution does.
Proposition 2.
@@
and with observations according to $\bm{y}=\bm{f_{0}}(\bm{X})+\bm{\varepsilon}$, we can calculate the expected squared differences between the estimated and true models, something that is often referred to as the risk: $\text{Risk}(\bm{\hat{\theta}};\bm{\theta_{0}})=\mathbb{E}\left(\|\bm{\hat{\theta}}-\bm{\theta_{0}}\|_{2}^{2}\right)$. In Proposition 3, which is an adaptation of Theorems 1 and 2 by Ali et al., (2019), the KGF estimation and prediction risks are bounded in terms of the corresponding KRR risks. In all three cases, the KGF risk is less than 1.69 times the KRR risk.
Proposition 3.
@@
To summarize this section, the KGF and KRR solutions tend to agree well both in terms of the parameter and prediction vectors and in terms of the risks. Which algorithm actually performs better, depends on the specific data, sometimes KGF performs slightly better than KRR and sometimes vice versa. Proposition 2 suggests that the KGF solution lies closer to the observations than KRR solution does, which can often be an advantage, however not in the presence of extreme outliers. These conclusions are supported by the experiments in Section 6.

### v7

@@ Abstract
Kernel ridge regression, KRR, is a generalization of linear ridge regression that is non-linear in the data, but linear in the model parameters. Here, we introduce an equivalent formulation of the objective function of KRR, which opens up for replacing the ridge penalty with the $\ell_{\infty}$ and $\ell_{1}$ penalties.
Using the $\ell_{\infty}$ and $\ell_{1}$ penalties, we obtain robust and sparse kernel regression, respectively. We study the similarities between explicitly regularized kernel regression and the solutions obtained by early stopping of iterative gradient-based methods, where we connect $\ell_{\infty}$ regularization to sign gradient descent, $\ell_{1}$ regularization to forward stagewise regression (also known as coordinate descent), and $\ell_{2}$ regularization to gradient descent, and, in the last case, theoretically bound for the differences. We exploit the close relations between $\ell_{\infty}$ regularization and sign gradient descent, and between $\ell_{1}$ regularization and coordinate descent to propose computationally efficient methods for robust and sparse kernel regression.
@@
Keywords: Robust Regression, Kernel Regression, Sign Gradient Descent, Gradient Descent, Gradient Flow
@@ 1 Introduction
In the linear case, many alternatives to the ridge penalty have been proposed, including the lasso penalty (Tibshirani,, 1996), which is known for creating sparse models. By replacing the ridge penalty of KRR with the lasso penalty Roth, (2004), Guigue et al., (2005) and Feng et al., (2016) have obtained non-linear regression that is sparse in the observations. Rather than using the lasso, or $\ell_{1}$, penalty, Russu et al., (2016) and Demontis et al., (2017) trained kernelized support vector machines with max (or infinity, $\ell_{\infty}$) norm regularization to obtain models that are robust against adversarial data, which in the context of regression translates to outliers.
@@
In this paper, we present an equivalent formulation of KRR, which we use to obtain kernel regression with the $\ell_{\infty}$ and $\ell_{1}$ norms, in addition to the default, $\ell_{2}$, norm. We also use the equivalent formulation to solve kernel regression using the three different gradient-based optimization algorithms gradient descent, sign gradient descent, and coordinate descent (forward stagewise regression). We relate each iterative algorithm to one explicitly regularized solution and use gradient-based optimization with early stopping to obtain computationally efficient robust, and sparse, kernel regression. Even though robust kernel regression has more evident applications than sparse kernel regression does, since the calculations are analogous for $\ell_{\infty}$ and $\ell_{1}$ regularization we include both cases. In the results section, we however focus on robust regression.
The rest of the paper is structured as follows. In Section 2, we review kernel ridge regression, before introducing our equivalent formulation in Section 3. In Section 4, we review kernel gradient descent, KGD, and present theoretical comparisons between KGD with infinitesimal step size and KRR. In Section 5, we generalize KRR into robust and sparse kernel regression by replacing the $\ell_{2}$ penalty with the $\ell_{\infty}$ and $\ell_{1}$ penalties, respectively. Similarly to in Section 4, we also introduce gradient-based optimization algorithms that we relate to the explicitly regularized methods; kernel sign gradient descent for robust kernel regression, and kernel coordinate descent for sparse kernel regression. Finally, in Section 6, we demonstrate our findings with experiments on real and synthetic data.
@@
All proofs are deferred to Appendix D.
@@ 2 Review of Kernel Ridge Regression
For a positive semi-definite kernel function, $k(\bm{x},\bm{x^{\prime}})\in\mathbb{R}$, and $n$ paired observations, $(\bm{x_{i}},y_{i})_{i=1}^{n}\in\mathbb{R}^{p}\times\mathbb{R}$, presented in a design matrix, $\bm{X}=[\bm{x_{1}},\bm{x_{2}},{\dots}\bm{x_{n}}]^{\top}\in\mathbb{R}^{n\times p}$, and a response vector, $\bm{y}\in\mathbb{R}^{n}$, and for a given regularization strength, $\lambda\geq 0$, the objective function of kernel ridge regression, KRR, is given by
@@
Here, $\bm{\hat{f}}=\bm{\hat{f}}(\bm{X})\in\mathbb{R}^{n}$ and $\bm{\hat{f}^{*}}=\bm{\hat{f}^{*}}(\bm{X^{*}},\bm{X})\in\mathbb{R}^{n^{*}}$ denote model predictions for the training data, $\bm{X}$, and new data, $\bm{X^{*}}=[\bm{x^{*}_{1}},\bm{x^{*}_{2}},{\dots}\bm{x^{*}_{n^{*}}}]^{\top}\in\mathbb{R}^{n^{*}\times p}$, and $\bm{K}=\bm{K}(\bm{X})\in\mathbb{R}^{n\times n}$ and $\bm{K^{*}}=\bm{K^{*}}(\bm{X^{*}},\bm{X})\in\mathbb{R}^{n^{*}\times n}$ denote two kernel matrices defined according to $\bm{K}_{ij}=k(\bm{x_{i}},\bm{x_{j}})$ and $\bm{K^{*}}_{ij}=k(\bm{x^{*}_{i}},\bm{x_{j}})$. The weighted norm, $\|\bm{v}\|_{\bm{A}}$, is defined according to $\|\bm{v}\|_{\bm{A}}^{2}=\bm{v}^{\top}\bm{A}\bm{v}$ for any symmetric positive definite matrix $\bm{A}$.
@@
An alternative interpretation of KRR is as linear regression for a non-linear feature expansion of $\bm{X}$. According to Mercer’s Theorem (Mercer,, 1909), every kernel can be written as the inner product of feature expansions of its two arguments: $k(\bm{x},\bm{x^{\prime}})=\bm{\varphi}(\bm{x})^{\top}\bm{\varphi}(\bm{x^{\prime}})$ for $\bm{\varphi}(\bm{x})\in\mathbb{R}^{q}$. Thus, denoting the feature expansions of the design matrix and the new data with $\bm{\Phi}=\bm{\Phi}(\bm{X})\in\mathbb{R}^{n\times q}$ and $\bm{\Phi^{*}}=\bm{\Phi^{*}}(\bm{X^{*}})\in\mathbb{R}^{n^{*}\times q}$, the two kernel matrices can be expressed as $\bm{K}=\bm{\Phi}\bm{\Phi}^{\top}$ and $\bm{K^{*}}=\bm{\Phi^{*}}\bm{\Phi}^{\top}$. For $\bm{\beta}=\bm{\Phi}^{\top}\bm{\alpha}$, Equations 1 and 3 become
@@ 3 Equivalent Formulations of Kernel Ridge Regression
In this section, we present an equivalent formulation of the objective function in Equation 1, which provides the same solution for $\bm{\hat{\alpha}}$. This formulation opens up for generalizing KRR by using penalties other than the ridge penalty and also provides an interesting connection to functional gradient descent (Mason et al.,, 1999).
In Appendix A, we take this one step further by reformulating Equation 1 directly in the model predictions, $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$, by presenting two equivalent objectives, that, when minimized with respect to $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$, generate the solution in Equation 3. In this and the following sections, all calculations are done with respect to $\bm{\alpha}$, and the corresponding expressions for $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ are obtained through multiplication with $[\bm{K}^{\top},\bm{K^{*}}{{}^{\top}}]^{\top}$, but, as shown in the appendix, the expressions for $[\bm{f}^{\top},\bm{f^{*}}{{}^{\top}}]^{\top}$ can be obtained directly without taking the detour over $\bm{\alpha}$.
In Proposition 3.1, we show how we can move the weighted norm in Equation 1 from the penalty term to the reconstruction term.
Proposition 3.1.
@@
Remark 3.1.
Despite the appearance of $\bm{K}^{-1}$ in Equation 5b, this inverse never needs to be calculated, since in all calculations, $\bm{K}^{-1}$ cancels through multiplication by $\bm{K}$.
The alternative formulation in Equation 5b, where the reconstruction term, rather than the regularization term, is weighted by the kernel matrix, has two interesting implications. First, a standard ridge penalty on the regularization term opens up for using other penalties than the $\ell_{2}$ norm, such as the $\ell_{1}$ and $\ell_{\infty}$ norms. Although these penalties have previously been used in the kernel setting, it has been done by replacing $\|\bm{\alpha}\|_{\bm{K}}^{2}$ in Equation 1 by $\|\bm{\alpha}\|_{1}$ or $\|\bm{\alpha}\|_{\infty}$, hence acting as if the objective of KRR were to minimize $\|\bm{y}-\bm{K}\bm{\alpha}\|_{2}^{2}+\lambda\|\bm{\alpha}\|_{2}^{2}$, and losing the connection to linear regression in feature space.
@@
$\bm{K}\bm{\alpha}-\bm{y}+\lambda\bm{\alpha}.$ (6)
@@
which is the gradient used in functional gradient descent. Functional gradient descent is generally derived by differentiating functionals. Here, however, the expression is a simple consequence of the equivalent objective function of KRR.
@@ 4 Kernel Gradient Descent and Kernel Gradient Flow
Before replacing the Euclidean norm in Equation 5b with the $\ell_{\infty}$ and $\ell_{1}$ norms in Section 5, we investigate solving kernel regression iteratively with gradient descent. We also use gradient descent with infinitesimal step size, known as gradient flow, to obtain a closed-form solution which we use for direct comparisons to kernel ridge regression.
The similarities between ridge regression and gradient descent with early stopping are well studied for linear regression (Friedman and Popescu,, 2004; Ali et al.,, 2019; Allerbo et al.,, 2023). When starting at zero, optimization time can be thought of as an inverse penalty, where longer optimization time corresponds to weaker regularization. When applying gradient descent to kernel regression, something we refer to as kernel gradient descent, KGD, we replace explicit regularization with implicit regularization through early stopping. That is, we use $\lambda=0$ and consider training time, $t$, as a regularizer.
@@
$\frac{d\bm{\hat{\alpha}}(t)}{dt}=\bm{y}-\bm{K}\bm{\hat{\alpha}}(t),\ \bm{\hat{\alpha}}(0)=\bm{0},$ (9)
the solution to which is given by Equation 10.
$\bm{\hat{\alpha}}_{\text{KGF}}(t)=(\bm{I}-\exp(-t\bm{K}))\bm{K}^{-1}\bm{y},$ (10)
@@
Remark 4.1.
Note that $(\bm{I}-\exp(-t\bm{K}))\bm{K}^{-1}=\bm{K}^{-1}(\bm{I}-\exp(-t\bm{K}))$ is well-defined even if $\bm{K}$ is singular. The matrix exponential is defined through its Taylor expansion and from $\bm{I}-\exp(-t\bm{K})=t\bm{K}-\frac{1}{2!}t^{2}\bm{K}^{2}+\dots$, a matrix $\bm{K}$ factors out, that cancels $\bm{K}^{-1}$.
Remark 4.2.
It is possible to generalize KGF to Nesterov accelerated gradient descent with momentum (Nesterov,, 1983; Polyak,, 1964). In this case $\exp(-t\bm{K})$ in Equation 10 generalizes to $\exp\left(-\frac{t}{1-\gamma}\bm{K}\right)$, where $\gamma\in[0,1)$ is the strength of the momentum. See Appendix D for details.
@@
where we have used that
$\left(\bm{K}+\lambda\bm{I}\right)^{-1}=\left(\bm{I}-\left(\bm{I}+1/\lambda\cdot\bm{K}\right)^{-1}\right)\bm{K}^{-1},$
which is a consequence of the Woodbury matrix identity.
@@ 4.1 Comparisons between Kernel Ridge Regression and Kernel Gradient Flow
Even though not explicitly stated, all estimates depend on the training data $(\bm{X},\bm{y})$. We further use the notation $\bm{\xi}\sim(\bm{\mu},\bm{\Sigma})$ do denote that the random variable $\bm{\xi}$ follows a distribution with $\mathbb{E}(\bm{\xi})=\bm{\mu}$ and Cov$(\bm{\xi})=\bm{\Sigma}$.
In Proposition 4.1, we bound the differences between the KGF and KRR solutions in terms of the non-regularized solutions, $\bm{\alpha_{0}}$, $\bm{\hat{f}}^{0}(\bm{X})$ and $\hat{f}^{0}(\bm{x^{*}})$. In parts (a) and (b), where we bound the differences between the parameter and in-sample prediction vectors, no further assumptions are needed. In parts (c) and (d), with bounds on individual parameters and predictions, including out-of-sample predictions, some very reasonable assumptions have to be made on the data. For all four bounds, the larger the non-regularized value is, the larger the difference between the KGF and KRR estimates is allowed to be.
Proposition 4.1.
@@
In Proposition 4.2, we compare the distances to the observation vector, $\bm{y}$, of the KGF and KRR solutions. According to parts (a) and (c), for a given regularization, the furthest possible (among all values of $\bm{y}$, or expected $y_{i}^{2}$) normalized KGF solution lies closer to the observations than the furthest possible normalized KRR solution does. Analogously, according to parts (b) and (d), for a given regularization, the furthest possible KRR solution lies closer to zero than the furthest possible KGF solution does. These results agree with those of Richards et al., (2021), who state that, for linear regression, when the signal-to-noise ratio is high, gradient descent tends to outperform ridge regression: For low noise, a solution that lies closer to the observations, as provided by KGF, is beneficial.
Proposition 4.2.
@@
and with observations according to $\bm{y}=\bm{f_{0}}(\bm{X})+\bm{\varepsilon}$, we can calculate the expected squared differences between the estimated and true models, something that is often referred to as the risk: $\text{Risk}(\bm{\hat{\theta}};\bm{\theta_{0}})=\mathbb{E}\left(\|\bm{\hat{\theta}}-\bm{\theta_{0}}\|_{2}^{2}\right)$. In Proposition 4.3, which is an adaptation of Theorems 1 and 2 by Ali et al., (2019), the KGF estimation and prediction risks are bounded in terms of the corresponding KRR risks. In all three cases, the KGF risk is less than 1.69 times the KRR risk.
Proposition 4.3.
@@
To summarize this section, the KGF and KRR solutions tend to agree well both in terms of the parameter and prediction vectors and in terms of the risks. Which algorithm actually performs better, depends on the specific data, sometimes KGF performs slightly better than KRR and sometimes vice versa. Proposition 4.2 suggests that the KGF solution lies closer to the observations than KRR solution does, which can often be an advantage, however not in the presence of extreme outliers. These conclusions are supported by the experiments in Section 6.